Solve the system of the following equations by either substitution or elimination:

3π‘₯π‘₯βˆ’2𝑦𝑦=14

4π‘₯π‘₯βˆ’3𝑦𝑦=17


1
Expert's answer
2021-06-16T12:26:31-0400

solving by elimination

"3xx-2yy=14...............................(1)\\4xx-3yy=17...............................(2)"


To eliminate xx multiply through eqn 1 by 4 and multiply through eqn 2 by 3


"12xx-8yy=56\\12xx-9yy=51"


Subtrating the second equation from the first equation of the previous system


"yy=5"


substituting the computed value in the equation 1

"3xx-2(5)=14
\\3xx-10=14\\3xx=14+10\\3xx=24\\xx=\frac{24}{3}\\xx=8"



xx=8

yy=5